If 3.00 g of magnesium hydroxide reacts according to the following equation how many mols of water. Will be produced?
Mg(oh)2 + 2hcl= 2h2o +mgo2
1
Expert's answer
2018-11-09T07:49:09-0500

M(Mg(OH)2) = 24+16*2+1*2 = 58g/mol

n(Mg(OH)2) = m/M = 3g/58g/mol = 0.052 mol

According to the equation n(H2O) = 0.052*2 = 0.104 mol